### v1

@@ 3.1 Overview
Our model architecture, schematically shown in Figure 4, is structured as a sequential combination of three primary components: i) a convolutional neural network (CNN) feature extractor, ii) a transformer encoder, and iii) a regression multi-layer perceptron (MLP) head. This hybrid architecture is the core strength of our model. The model combines a CNN feature extractor ($\mathcal{F}$) with a Transformer encoder ($\mathcal{T}$), leveraging the complementary strengths of both architectures for analysing time-series data [foumani2023deeplearningtimeseries]. Whereas CNNs are highly effective at efficiently identifying local patterns and hierarchical structures within the input signal, the Transformer’s self-attention mechanism excels at modeling global context and capturing subtle long-range temporal dependencies.
@@
    Transformer Encoder (T): The latent sequence of feature vectors ($L$) is first linearly projected to match the transformer’s input dimensionality, to be followed by the addition of positional encodings to inject information about the position of each element in the sequence. To facilitate sequence-level aggregation, a special token, often denoted as [CLS], is also prepended to the sequence before the addition of positional encodings. This approach, popularized by models like BERT [bert], allows the model to learn an aggregate representation of the entire input sequence into the output state corresponding to this [CLS] token. The encoder then processes this augmented sequence ($L^{\prime}$) using stacked layers of self-attention and feed-forward networks to capture global dependencies and contextual relationships within the data, outputting an encoded sequence representation ($E$). We write,
@@ 3.2 Feature extractor
The CNN feature extractor ($\mathcal{F}$), illustrated in Figure 2, is based on the InceptionTime architecture [inceptiontime], originally proposed for time-series classification. The network begins with a dedicated (stem) layer which processes the raw input signal ($\mathcal{X}$), which accepts a single-channel time series, typically of 4096 samples, using a convolution with a kernel size $k=7$ and a stride equal to 1. This convolution is followed by Batch Normalization (BN) [ioffe2015batchnormalizationacceleratingdeep] and a ReLU activation. The stem layer serves as a preparatory phase, transforming the input into an initial, richer feature representation, by projecting the input features to a higher dimensionality – we use $64$ channels as default. The core of the feature extractor is constructed from a sequence of $N_{B}$ main processing blocks stacked sequentially. Each of these main blocks internally performs a sequence of operations. First, the block processes its input through a stack of $M$ InceptionModule sub-blocks, illustrated in Figure 3. A residual connection [deepresiduallearning] is then applied around this entire stack of $M$ modules, and the result is passed through a ReLU activation. Finally, each main block is concluded by a MaxPool layer for temporal downsampling, with a kernel size and stride of 2. The latter operation reduces the sequence length (halving it at each step) and the computational cost for subsequent layers, which enables the model to learn features at broader time scales. Unlike the original InceptionTime architecture which uses Global Average Pooling for a final classification, our architecture employs this progressive MaxPool downsampling between blocks. This is an important adaptation, motivated by the computational complexity of the subsequent Transformer encoder. The self-attention mechanism in a standard Transformer has a computational and memory complexity of $O(N^{2})$, where $N$ is the sequence length. This quadratic scaling makes processing very long sequences, such as our raw GW time series, computationally prohibitive. By progressively downsampling with a stride of 2, we halve the sequence length at each step, and so reducing the computational load of the attention mechanism by a factor of four. This makes the application of a Transformer encoder more feasible while still preserving the sequential nature of the data and enabling hierarchical feature learning. While we use the standard self-attention, we note that significant research has been dedicated to mitigating this quadratic bottleneck, leading to more efficient Transformer variants with linear-time attention mechanisms [tay2022efficienttransformerssurvey].
@@
The InceptionModule follows the core idea of the Inception design, that is, parallel convolutional branches processing the input at multiple scales, with outputs concatenated along the channel dimension. Our implementation adapts the original InceptionTime formulation [inceptiontime] with a key modification to its normalization strategy. Instead of applying Batch Normalization after each internal convolutional layer, we apply a single Batch Normalization layer only after the outputs of all parallel branches have been concatenated. This design choice aims to reduce the overall computational cost of the module and provides a unified normalization step on the aggregated multi-scale feature map, which can help stabilise training by controlling the distribution of the combined features before they are passed to subsequent layers.
Following the original idea of InceptionTime block architecture [inceptiontime], we use a bottleneck layer based on a convolution with kernel $k=1$ and stride 1, which projects the channel dimension of the input time series to a smaller number. We set the default bottleneck size value to 32 output channels. This bottleneck layer significantly reduces dimensionality and model complexity, allowing for the use of larger effective filters in subsequent layers with a comparable parameter count. A ReLU activation is then applied after the bottleneck operation. The output of this path is then processed by three parallel convolutional layers with a stride of 1 and with different kernel sizes, enabling the capture of temporal patterns at different scales. For the size of the three kernels we choose $k\in\{10,20,40\}$. The use of such large kernel sizes is a key feature of the InceptionTime design, made computationally feasible by the preceding bottleneck layer. These large kernels allow the model to capture features over extended time horizons, which is particularly suitable for the potentially long-duration and complex morphologies of GW signals. We also use padding to preserve the sequence length through the module, so the output of all convolutional branches has the same sequence length. Each convolution is configured to output a common number of feature maps, $d_{branch}$, which is constant within a main block but can vary between main blocks. We moderately increase the number of filters ($d_{\text{branch}}$) compared to the design of Ref. [inceptiontime] in deeper blocks of the network. This design compensates for the reduction in temporal resolution caused by the MaxPool downsampling and allows the network to learn progressively more complex feature representations from the compressed sequence. Concurrently, a separate MaxPool pathway processes the original input ($x$), with $k=3$ and stride 1, which helps the model achieve some invariance to small temporal perturbations [inceptiontime]. The output of this MaxPooling operation is then passed through a convolution layer with kernel size $k=1$ to project its feature dimension to match the filters of the convolutional branches.
Finally, the outputs from all three parallel convolutional branches and the processed MaxPooling pathway are concatenated along their feature channel dimension and passed through a final BatchNorm (BN) layer. The entire feature extractor network leads to a final Batch Normalization layer which normalises the output feature sequence $L$. For all convolutional layers, if pretrained weights are not used, the weights are initialised using Kaiming Normal initialization [kaiming], which is well-suited for the ReLU activations used throughout the network. In our experiments, the feature extractor was configured with $N_{B}=3$ main blocks, each containing $M=3$ InceptionModules. The bottleneck size within each module was set to $d_{bottle}=32$. The number of filters per branch, $d_{branch}$, was varied for each of the main blocks, with values of $[32,48,64]$ respectively.
@@ 3.3 Transformer encoder and regression head
The resulting latent sequence of the feature extractor ($\mathcal{F}$) has dimensions $L\in\mathbb{R}^{B\times N\times d_{\mathcal{F}}}$, where $B$ is the batch size, $N$ is the sequence length reduced by $\mathcal{F}$, and $d_{\mathcal{F}}$ is the feature dimension (output channels) from $\mathcal{F}$. This sequence is first projected into the transformer’s model dimension ($d_{model}$). We use a linear projection layer with weight matrix $W_{0}\in\mathbb{R}^{d_{\mathcal{F}}\times d_{model}}$ and bias $b_{0}\in\mathbb{R}^{d_{model}}$, applied independently to each element along the sequence length dimension for all batches as
$L_{(b,i,\cdot)}=L_{(b,i,\cdot)}W_{0}+b_{0},\quad\text{for }b=1..B,i=1..N,$ (10)
which leads to the projected sequence with dimensions $L\in\mathbb{R}^{B\times N\times d_{model}}$. As described previously, a special classification token, [CLS], is prepended to $L$ forming the input sequence $L^{\prime}$, allowing the model to learn a sequence-level summary through the self-attention mechanism [bert]. Positional encodings are added to $L^{\prime}$ to provide the model with sequence order information, which the transformer itself does not inherently capture. We use standard sinusoidal encoding as in [vaswani2017attention].
Our encoder follows the standard transformer architecture [vaswani2017attention], and consists of a stack of identical layers, each comprising a multi-head self-attention mechanism and position-wise fully connected feed-forward network. Residual connections are employed around each of the two sub-layers, followed by layer normalization. GELU activation is used. The encoder processes the entire augmented sequence $L^{\prime}$ and outputs an encoded sequence $E$ of the same length, where each output vector represents its corresponding input element enriched with contextual information from the entire sequence. The output vector corresponding to the [CLS] token acts as an aggregated representation $E_{agg}$ of the whole input sequence which is then fed into the MLP regression head $\mathcal{H}$, consisting of three fully-connected linear layers, followed by GELU and dropout, except for the final output layer. $\mathcal{H}$ maps the learned, feature representation $E_{agg}$ to the inferred parameters, $y=\mathcal{H}(E_{agg})$.
@@ 4.1 Data pre-processing
The entire dataset was initially shuffled randomly to ensure an unbiased distribution of samples, and then partitioned into a training set ($D_{train}$), a validation set ($D_{val}$), and a test set ($D_{test}$) under a split ratio of 8080%, 1010% and 1010% respectively. $D_{train}$ was used for learning the model parameters $\theta$, $D_{val}$ was used to monitor the model’s performance on data not seen during training, to guide the selection of the best-performing model checkpoint, and to inform the hyperparameter optimization process. $D_{test}$ was held out during all training and model selection stages and was used at the end to provide an unbiased estimate of the final selected model’s generalization performance, including the test error and other relevant evaluation metrics.
@@ 4.2 Pretraining and training
We adopt a self-supervised pretraining strategy inspired by pretraining task proposed in [zerveas2020transformerbasedframeworkmultivariatetime]. This self-supervised objective is conceptually similar to the Masked Language Model (MLM) pretraining task popularized by BERT [bert], in that the model learns contextual representations by reconstructing a corrupted input. However, our implementation is adapted for the continuous time-series domain and differs in several key aspects. First, while BERT operates on a sequence of discrete text tokens from a finite vocabulary, our model processes a continuous-valued signal. Second, our masking strategy differs: BERT typically replaces a small percentage (usually around 15%) of individual random tokens with a special ‘[MASK]‘ token, whereas we mask contiguous spans of the time series by setting their values to zero. This span-based masking encourages the model to learn the structure of longer temporal segments. Finally, the prediction tasks are distinct: BERT performs a classification task, predicting the original discrete token ID using a Cross-Entropy loss. In contrast, our model performs a regression task, predicting a continuous latent vector ($\mathbf{z}_{\text{pred}}$) and is optimised using a MSE loss against the target latent vector ($\mathbf{z}_{\text{target}}$) derived from the uncorrupted signal. During pretraining, random contiguous segments of input waveforms are masked (set to zero), and the model is tasked with reconstructing a latent representation of the original, uncorrupted signal from this partial input. This process encourages the model to learn robust temporal dependencies and feature representations. Given an input waveform $\mathbf{x}\in\mathbb{R}^{L}$, the binary mask $\mathbf{m}\in\{0,1\}^{L_{in}}$ is generated stochastically for each sample using a process modeled as a two-state discrete-time Markov chain (see Figure 5). The two states are $S_{0}$ (being inside a masked segment, where the mask value is 0) and $S_{1}$ (being inside an unmasked segment, where the mask value is 1). The duration of stay in each state follows a geometric distribution. The probability of transitioning out of a state at any given time step is denoted by $p$. The mean segment length is therefore $1/p$. We define the mean length for masked segments as $l_{m}$ and for unmasked segments as $l_{u}$, which in turn define the transition probabilities $p_{m}=1/l_{m}$ and $p_{u}=1/l_{u}$, respectively. The process is determined by two key hyperparameters: the desired mean masked length, $l_{m}$, and the overall masking ratio, $r$, from which the mean unmasked length is derived as $l_{u}=l_{m}\cdot(1-r)/r$. The transition matrix $\mathbf{P}$ for this Markov chain is given by
@@
For our experiments, we set a masking ratio of $r=0.15$ and a mean masked segment length of $l_{m}=32$. This choice of a relatively long mean segment length is deliberate and tied to our network’s architecture. It serves two purposes: first, it prevents the reconstruction task from becoming trivial, as the model must learn long-range dependencies rather than simply interpolating from immediate neighbors. Second, and more critically, it ensures that the masked regions are robust to the temporal downsampling performed by the feature extractor. With a total downsampling factor of 8x in our network, a mean masked length of 32 ensures that each masked segment in the input corresponds to multiple tokens in the latent space where the reconstruction loss is computed, providing a stable and consistent learning signal. The masked vector is then applied to the input waveform via element-wise multiplication to produce the masked input
@@
To prevent the self-attention mechanism from attending to padded regions, the original padding mask $\mathbf{p}$ is downsampled to the new sequence length $L^{\prime}$. This downsampling must precisely mirror the temporal reduction performed by the main feature extractor $\mathcal{F}$. We achieve this by applying a sequence of MaxPool layers (with a kernel size and stride of 2, identical to the downsamplers in $\mathcal{F}$) to the padding mask. Using Max Pooling on the binary mask correctly propagates the padding information, ensuring that a position in the resulting downsampled mask, $\mathbf{p}^{\prime}\in\{0,1\}^{L^{\prime}}$, is marked as padding (value = 0) only if all input positions in its receptive field were also padding. The resulting mask is then inverted to define the attention mask $\mathbf{m}_{\text{attn}}=\neg\mathbf{p}^{\prime}$, which is supplied to the transformer encoder. The transformer encoder $\mathcal{T}(\cdot)$ processes the input and the output is passed through a linear prediction head $f_{\text{head}}(\cdot)$ as
@@
Following pretraining, the model is trained/fine-tuned to predict the target physical parameters $\mathbf{y}_{i}$ (e.g masses, spins) using supervised training. We minimize the MSE over batches $\mathcal{B}\subset D_{\text{train}}$,
@@
where $f(\mathbf{x}_{i};\theta)$ is the predicted target vector and $\theta$ the deep model’s inherent parameters. Fine-tuning uses the AdamW optimiser [loshchilov2019decoupledweightdecayregularization] with learning rate $5e^{-4}$, cosine scheduler with linear 10% warmup, batch size $B=250$, and a maximum of $E=120$ epochs. Gradient clipping with threshold 1.0 is applied to prevent instability. The hyperparameters of our model are summarized in Table 10. We evaluate the trained model, which we denote as $f(\mathbf{x};\theta^{**})$, on the test dataset ($D_{\text{test}}$) using two regression metrics for each predicted physical parameter. Let us denote with $y_{ij}$ and $\hat{y}_{ij}$ the true and predicted values of the $j$-th parameter for the $i$-th test sample after inverse scaling, and with $\bar{y}_{j}$ the mean of the true $j$-th parameter over $D_{\text{test}}$ respectively. We then define the metrics
@@
The Mean Absolute Error (MAE) captures average absolute errors, while $R^{2}_{j}$ measures the proportion of variance explained by the model. Perfect predictions yield $R^{2}_{j}=1$, while $R^{2}_{j}<0$ indicates worse performance than predicting the mean.

### v2

@@ 3.1 Overview
Our model architecture, schematically shown in Figure 4, is structured as a sequential combination of three primary components: i) a convolutional neural network (CNN) feature extractor, ii) a transformer encoder, and iii) a regression multi-layer perceptron (MLP) head. This hybrid architecture is the core strength of our model. The model combines a CNN feature extractor ($\mathcal{F}$) with a Transformer encoder ($\mathcal{T}$), leveraging the complementary strengths of both architectures for analysing time-series data. Whereas CNNs are highly effective at efficiently identifying local patterns and hierarchical structures within the input signal, the Transformer’s self-attention mechanism excels at modelling global context and capturing subtle long-range temporal dependencies.
@@
    Transformer Encoder (T): The latent sequence of feature vectors ($L$) is first linearly projected to match the transformer’s input dimensionality, to be followed by the addition of positional encodings to inject information about the position of each element in the sequence. To facilitate sequence-level aggregation, a special token, often denoted as [CLS], is also prepended to the sequence before the addition of positional encodings. This approach, popularised by models like BERT [23], allows the model to learn an aggregate representation of the entire input sequence into the output state corresponding to this [CLS] token. The encoder then processes this augmented sequence ($L^{\prime}$) using stacked layers of self-attention and feed-forward networks to capture global dependencies and contextual relationships within the data, outputting an encoded sequence representation ($E$). We write,
@@ 3.2 Feature extractor
The CNN feature extractor ($\mathcal{F}$), illustrated in Figures 2 and 3, is based on the InceptionTime architecture [24], adapted for our hybrid pipeline. A stem layer (convolution with $k=7$, followed by BN and ReLU) projects the single-channel input to 64 feature channels. The core consists of $N_{B}=3$ main blocks, each containing a stack of $M=3$ InceptionModule sub-blocks with a residual connection, followed by a MaxPool layer (kernel size and stride of 2) for temporal downsampling. This progressive downsampling is a departure from the original InceptionTime design, which uses Global Average Pooling for classification. Here, halving the sequence length at each block reduces the quadratic cost of the subsequent Transformer’s self-attention calculation by a factor of four per block, making the hybrid architecture more computationally feasible for potentially long GW time series. While we use the standard self-attention, we note that significant research has been dedicated to mitigating this quadratic bottleneck, leading to more efficient Transformer variants with linear-time attention mechanisms [44]. Each InceptionModule uses a bottleneck ($k=1$, $d_{\text{bottle}}=32$) followed by three parallel convolutions with kernels $k\in\{10,20,40\}$ alongside a MaxPool pathway ($k=3$, stride 1, preserving sequence length) for local shift invariance. We make two modifications relative to the original InceptionTime formulation [24]: (i) we apply a single Batch Normalization layer after concatenating all branch outputs, rather than after each internal convolution, reducing computational cost while providing a unified normalization of the multi-scale features; and (ii) we progressively increase the number of filters per branch across blocks, $d_{\text{branch}}\in\{32,48,64\}$, to compensate for the temporal compression introduced by MaxPool downsampling.
A convolutional kernel of size $k$ at sampling rate $f_{s}$ spans a time window $\Delta t=k/f_{s}$ and can resolve features with frequencies approximately $f\gtrsim f_{s}/k$. For GW signals, whose frequency increases from inspiral to merger, the branches with $k\in\{10,20,40\}$ act as complementary band-pass filters: larger kernels have the capacity to capture the slowly varying inspiral chirp (chirp-mass dominated), while smaller kernels resolve the rapid merger–ringdown oscillations encoding mass ratio and spin. This multi-scale design matches the chirping structure of compact binaries. The kernel choices were empirically motivated. The total downsampling $2^{N_{B}}=8$ yields a latent resolution of $\sim$2 ms per token, adequate for our mass range. Future extensions could employ adaptive kernels and learnable downsampling. The fixed sampling rate of 4096 Hz and input length of $4096$ samples limit generality. Future improvements include (i) multi-rate training with augmentation to promote rate-invariant representations, and (ii) an input resampling layer that maps arbitrary sampling to a standardised representation with adjusted positional encodings.
@@ 3.3 Transformer encoder and regression head
The feature extractor produces a latent sequence $L\in\mathbb{R}^{B\times N\times d_{\mathcal{F}}}$, which is linearly projected to the transformer’s model dimension $d_{\text{model}}=256$. A learnable [CLS] token [23] is prepended and standard sinusoidal positional encodings [48] are added, forming the input sequence $L^{\prime}$. Our encoder follows the standard Transformer architecture [48] with 4 layers, 4 attention heads, feed-forward dimension $d_{ff}=512$, GELU activations, and a dropout rate of 0.10. The output vector at the [CLS] position serves as the sequence-level representation $E_{\text{agg}}$, which is passed to the regression MLP head $\mathcal{H}$: two hidden layers (dimensions 256 and 64) with GELU and dropout, followed by a linear output layer mapping to the four inferred parameters, $\mathbf{y}=\mathcal{H}(E_{\text{agg}})$.
@@ 4.1 Data pre-processing
The entire dataset was initially shuffled randomly to ensure an unbiased distribution of samples, and then partitioned into a training set ($D_{train}$), a validation set ($D_{val}$), and a test set ($D_{test}$) under a split ratio of 8080%, 1010% and 1010% respectively. $D_{train}$ was used for learning the model parameters $\theta$, $D_{val}$ was used to monitor the model’s performance on data not seen during training and to guide the selection of the best-performing model checkpoint. $D_{test}$ was held out during all training and model selection stages and was used at the end to provide an unbiased estimate of the final selected model’s generalisation performance, including the test error and other relevant evaluation metrics.
@@ 4.2 Pretraining and training
We adopt a self-supervised pretraining strategy inspired by pretraining task proposed in [54]. This self-supervised objective is conceptually similar to the Masked Language Model (MLM) pretraining task popularised by BERT [23], in that the model learns contextual representations by reconstructing a corrupted input. However, our implementation is adapted for the continuous time-series domain and differs in several key aspects. First, while BERT operates on a sequence of discrete text tokens from a finite vocabulary, our model processes a continuous-valued signal. Second, our masking strategy differs: BERT typically replaces a small percentage (usually around 15%) of individual random tokens with a special ‘[MASK]‘ token, whereas we mask contiguous spans of the time series by setting their values to zero. This span-based masking encourages the model to learn the structure of longer temporal segments. Finally, the prediction tasks are distinct: BERT performs a classification task, predicting the original discrete token ID using a Cross-Entropy loss. In contrast, our model performs a regression task, predicting a continuous latent vector ($\mathbf{z}_{\text{pred}}$) and is optimised using a MSE loss against the target latent vector ($\mathbf{z}_{\text{target}}$) derived from the uncorrupted signal. During pretraining, random contiguous segments of input signals are masked (set to zero), and the model is tasked with reconstructing a latent representation of the original, uncorrupted signal from this partial input. This process encourages the model to learn temporal dependencies and feature representations. Given an input waveform $\mathbf{x}\in\mathbb{R}^{L}$, the binary mask $\mathbf{m}\in\{0,1\}^{L_{in}}$ is generated stochastically for each sample using a process modeled as like a two-state discrete-time Markov chain (see Figure 5). The two states are $S_{0}$ (being inside a masked segment, where the mask value is 0) and $S_{1}$ (being inside an unmasked segment, where the mask value is 1). The duration of stay in each state follows a geometric distribution. The probability of transitioning out of a state at any given time step is denoted by $p$. The mean segment length is therefore $1/p$. We define the mean length for masked segments as $l_{m}$ and for unmasked segments as $l_{u}$, which in turn define the transition probabilities $p_{m}=1/l_{m}$ and $p_{u}=1/l_{u}$, respectively. The process is determined by two key hyperparameters: the desired mean masked length, $l_{m}$, and the overall masking ratio, $r$, from which the mean unmasked length is derived as $l_{u}=l_{m}\cdot(1-r)/r$. The transition matrix $\mathbf{P}$ for this Markov chain is given by
@@
In practice, the implementation uses strictly alternating masked and unmasked segments, which is a close approximation to the Markov chain for the chosen parameters. For our experiments, we set a masking ratio of $r=0.15$ and a mean masked segment length of $l_{m}=32$. This choice of a relatively long mean segment length is tied to our network’s architecture. It serves two purposes: first, it prevents the reconstruction task from becoming trivial, as the model must learn long-range dependencies rather than simply interpolating from immediate neighbors. Second, and more critically, it ensures that the masked regions are robust to the temporal downsampling performed by the feature extractor. With a total downsampling factor of 8x in our network, a mean masked length of 32 ensures that each masked segment in the input corresponds to multiple tokens in the latent space where the reconstruction loss is computed, providing a stable and consistent learning signal. The masked vector is then applied to the input waveform via element-wise multiplication to produce the masked input
@@
To prevent the self-attention mechanism from attending to padded regions, the original padding mask $\mathbf{p}$ is downsampled to the new sequence length $L^{\prime}$. This downsampling must precisely mirror the reduction performed by the main feature extractor $\mathcal{F}$. We achieve this by applying a sequence of MaxPool layers (with a kernel size and stride of 2, identical to the downsamplers in $\mathcal{F}$) to the padding mask. Using Max Pooling on the binary mask correctly propagates the padding information, ensuring that a position in the resulting downsampled mask, $\mathbf{p}^{\prime}\in\{0,1\}^{L^{\prime}}$, is marked as padding (value = 0) only if all input positions in its receptive field were also padding. The resulting mask is then inverted to define the attention mask $\mathbf{m}_{\text{attn}}=\neg\mathbf{p}^{\prime}$, which is supplied to the transformer encoder. The transformer encoder $\mathcal{T}(\cdot)$ processes the input and the output is passed through a linear prediction head $f_{\text{head}}(\cdot)$ as
@@
The evolution of the pretraining loss is shown in Figure 1(a) in the Appendix. We select the pretrained weights at the epoch of minimum validation loss (denoted with a star in Figure 1(a)). The downstream improvements reported in Table 2 confirm that the representations learned during the convergence phase are effective.
Following pretraining, the model is trained/fine-tuned to predict the target physical parameters $\mathbf{y}_{i}$ (e.g masses, spins) using supervised training. We minimise the MSE over batches $\mathcal{B}\subset D_{\text{train}}$,
@@
where $f(\mathbf{x}_{i};\theta)$ is the predicted target vector and $\theta$ the deep model’s inherent parameters. Fine-tuning uses the AdamW optimiser [35] with learning rate $5\cdot 10^{-4}$, cosine scheduler with linear 10% warmup, batch size $B=250$, and a maximum of $E=120$ epochs. Gradient clipping with threshold 1.0 is applied to prevent instability. The hyperparameters of our model are summarised in Table A1. We evaluate the trained model, which we denote as $f(\mathbf{x};\theta^{**})$, on the test dataset ($D_{\text{test}}$) using two regression metrics for each predicted physical parameter. Let us denote with $y_{ij}$ and $\hat{y}_{ij}$ the true and predicted values of the $j$-th parameter for the $i$-th test sample after inverse scaling, and with $\bar{y}_{j}$ the mean of the true $j$-th parameter over $D_{\text{test}}$ respectively. We then define the metrics
@@
The Mean Absolute Error (MAE) captures average absolute errors, while $R^{2}_{j}$ measures the proportion of variance explained by the model. Perfect predictions yield $R^{2}_{j}=1$, while $R^{2}_{j}<0$ indicates worse performance than predicting the mean. We note that R² is well-defined regardless of the distribution of target values. On our uniform grid, it provides a conservative assessment since the model is evaluated with equal weight across the full parameter range rather than being dominated by the most probable astrophysical configurations.